A study was made by a publishing company on how advertising costs for the company correlated with sales.

  1. Find the equation of the regression line to predict sales from advertising costs.
  2. Estimate the weekly sales when advertising costs are 35TL.

Advertising costs

Sales

50

530

35

410

45

360

90

790

60

710

60

480
Answer #1

х Y X*Y x2 y2 50 530 26500 2500 280900 35 410 14350 1225 168100 45 360 16200 2025 129600 90 790 71100 8100 624100 60 710 4260from above table we calculate340 X = X; = 56.666 6 η 3280 Y Σ = 546.666 n 6 i=1 2 SSxx = ΣΧ? Σx: (Σ = 21050 - 340? /6 = 1783.333 i=1 2 (Σχ.) 1850, - Σκι -

Let m be the slope and n be the y intercept

SSXY m = 13683.333 1783.333 = 7.6729 SSxx n= Ï - Ì. . m = 546.666 - 56.666x7.6729 = 111.8692 regression equation is: Y = 111.according to question when X is 35

So according to regression equation sales for 35TL is

Y=111.8692+7.6729×35= 380.4207